OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

Will Fan and Medal! Thanks!

OpenStudy (anonymous):

Please help! @gerryliyana

OpenStudy (anonymous):

I think so far that the answer is C

OpenStudy (anonymous):

check first 6x + 2y = -6 | times 2| 3x - 4y = -18 then 12x + 4y = -12 3x - 4y = -18 --------------------+ 15x = -30

OpenStudy (anonymous):

Thank you! Could you help me with one more?

OpenStudy (anonymous):

of course go ahead

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

my computer is going crazy

OpenStudy (anonymous):

ok i'll wait you

OpenStudy (anonymous):

Solve the system of equations. 2x - 5y = 3 x - 3y = 1 (4, 1) (5, 2) (7, 2) (9, 3)

OpenStudy (anonymous):

I think it is D

OpenStudy (anonymous):

2x - 5y = 3 x - 3y = 1 |times 2| then 2x - 5y = 3 2x - 6y = 2 --------------- - y = 1 then check fo x by using x - 3y = 1 x - 3y = 1 x = 1 + 3y x = 1 + 3(1) x = 4

OpenStudy (anonymous):

So it should be A

OpenStudy (anonymous):

Oh, ok.. I get where I went wrong.. Once again, thank you soo much!

OpenStudy (anonymous):

you are very welcome @sara_lee101